# Training Large Language Models to Reason In Parallel with Global Forking Tokens

### v2

@@ Abstract
Although LLMs have demonstrated improved performance by scaling parallel test-time compute, doing so relies on generating reasoning paths that are both diverse and accurate. For challenging problems, the forking tokens that trigger diverse yet correct reasoning modes are typically deep in the sampling tree. Consequently, common strategies to encourage diversity, such as temperature scaling, encounter a worsened trade-off between diversity and accuracy. Motivated by this challenge, we treat parallel reasoning as a set-of-next-token-prediction problem, and incorporate a set-based global loss into Supervised Fine-Tuning (SFT) using self-supervised bipartite matching between our global forking tokens and unique reasoning traces. We observe that, while naive fine-tuning with multiple reasoning traces collapses these unique reasoning modes, our proposed method, Set Supervised Fine-Tuning (SSFT), preserves these modes and produces emergent global forking tokens. Experiments on multiple reasoning benchmarks show that our SSFT consistently outperforms SFT under both $\mathrm{Pass@}1$ and $\mathrm{Cons@}k$ metrics.
Figure 1: An illustration of different supervised fine-tuning methods that aim to instill parallel reasoning capabilities from diverse reasoning traces for the same question. Compared to (1) standard SFT and (2) SFT with randomly assigned parallel thinking identifiers, (3) Set-Supervised Fine-Tuning leverages a self-supervised bipartite matching process to learn to maximally differentiate the reasoning modes conditioned on distinct <think i>, for each question. Self-supervised matching prevents collapse of maximally distinct reasoning modes caused by ordering bias from randomly or manually assigning a reasoning mode identifier to a reasoning trace. The learned global forking tokens are considered emergent because we observe that similar reasoning modes over the traces for different questions cluster to match the same <think,i> at convergence, even though we do not manually identify such modes or incorporate any regularization term to maintain consistency in matching over different input prompts.
@@ 1 Introduction
Large language models have recently improved reasoning by allocating more test-time compute to generate more tokens before producing the final answer (OpenAI2025mini). However, extended sequential scaling can lead to “overthinking”, where performance decreases beyond a certain sequence length (ghosal2025doesoverthinkinghelp; chen2024notoverthinkthatmuch). To mitigate this, another scaling dimension based on repeated parallel sampling and aggregation (wang2022selfconsistency; brown2024large_repeatedsampling) has shown success in further boosting reasoning performance. However, these methods rely on LLMs generating diverse yet correct solutions; as tasks become harder, a new mechanism is required to increase diversity. Recent work shows that only a minority of tokens in Chain-of-Thought reasoning (wei2022chain) can act as forking tokens that lead to distinct reasoning modes (wang2025beyond_80_20_rule; wang2025emergenthierarchicalreasoning), so as the problem becomes harder and the generation becomes longer, it can become substantially harder to sample them. Also, common practices to encourage diversity, typically through temperature scaling, inherently entail a diversity-accuracy trade-off, as the forking tokens that trigger diverse yet correct reasoning modes are typically located deeply within the sampling tree. Moreover, recent theoretical work also shows that increasing the temperature alone does not necessarily guarantee greater diversity unless the model is explicitly trained for coverage. (verine2025improvingwhentemperaturefailschangetheloss).
@@
Building on these observations, we aim to leverage diverse reasoning traces to train for coverage (guo2025deepseekR1; gemini25). We introduce global forking tokens to initiate parallel reasoning traces and frame parallel reasoning as a set prediction problem. Specifically, given a question, an LLM, conditioned on a reserved set of tokens in a chosen ordering, generates $\mathrm{M}$ reasoning sequences in parallel, each aligned with one of $\mathrm{M}$ ground-truth reasoning traces. For each ordering, we compute the total autoregressive loss across the generated sequences. By enumerating all possible orderings, we identify the minimum loss, which defines the set language modeling loss, conditioned on the distinct forking tokens (Equation 5). This formulation naturally incorporates coverage into the training objective and is capable of learning global forking tokens that can serve as prompts to trigger reasoning modes that are both diverse and accurate. We operationalize this idea through our Set Supervised Fine-Tuning (SSFT) framework. Our main contributions are summarized below.
@@
  • We introduce global forking tokens and incorporate a set-based loss into SFT using bipartite matching between a set of reserved global forking tokens and diverse reasoning traces (Section 2; Figure 2). The goal is that, after training with SSFT, the model can initiate distinct reasoning modes when prompted with different tokens from this set, thus reducing the dependence on sampling forking tokens mid-generation (Section 2.3; Figure 3).
  • We empirically demonstrate that, across multiple reasoning benchmarks, a model fine-tuned with SSFT outperforms a standard SFT model trained on the same distilled traces from diverse teacher models, improving $\mathrm{Pass@}1$, $\mathrm{Pass@}k$, and $\mathrm{Cons@}k$ (Section 3, Table 1 and Figure 4). We also find that naively fine-tuning using diverse reasoning traces without the set loss from bipartite matching can result in such control tokens initiating the same collapsed reasoning mode (Figure 6). In contrast, the global forking tokens learned by SSFT maximally differentiate reasoning modes (Figure 5).
  • To facilitate training where different input prompts can use different numbers of parallel reasoning traces, which can cause difficulty in using gradient accumulation under distributed training, we also include a simple implementation based on the HuggingFace repository. Our implementation does not concatenate a variable number of reasoning traces along the sequence dimension and modify causal attention, which would require sequence parallelism and more complicated fused kernels to avoid out-of-memory for training with long reasoning traces (Appendix A.1).
@@ 2 Learning Global Forking Tokens via Set Supervised Finetuning
Background on Language Modeling with Reasoning. In language modeling, the goal is to train a model $\pi_{\bm{\theta}}$ to approximate the joint distribution over a sequence of word tokens ${\mathbf{x}}=\{x_{i}\}_{i=1}^{\mathrm{T}}\in\mathcal{V}^{\mathrm{T}}$, where $\mathrm{T}$ is the sequence length, and each token is within a finite vocabulary set $\mathcal{V}$. An autoregressive model uses the chain rule to represent it as a product of conditionals on the preceding tokens:
$\pi_{\bm{\theta}}({\mathbf{x}})=\prod_{t=1}^{\mathrm{T}}\pi_{\bm{\theta}}\left({\textnormal{x}}_{t}|{\mathbf{x}}_{<t}\right)$ (1)
[…]
@@ 2.1 Parallel Reasoning as Set of Next Token Prediction
Problem Setup. In this paper, our goal is not only to instill new reasoning capabilities into a model, but also to ensure that prompting with a set of reserved special tokens, in parallel with a question, elicits distinct reasoning traces. We call these global forking tokens ${\bm{g}}:=\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ instantiated as $\{\texttt{<think\,i>\,}\}_{i=1}^{\mathrm{N}}$ tags. We use ${\textnormal{g}}^{(i)}$ interchangeably with <thinki>, depending on context for clarity. We consider a setting with multiple sources of reasoning traces, obtained at low cost without human annotation by distilling from diverse teachers, sampling repeatedly, and potentially using rejection sampling to filter out low-quality or incorrect traces. We adopt this low-cost regime to highlight the effectiveness of our algorithm, though the method extends naturally to settings with well-annotated, human-labeled data. So our problem is to (1) learn to do a set of next-token predictions on multiple distinct yet correct reasoning traces ${\mathbf{R}}:=\{{\mathbf{r}}^{(j)}\}_{j=1}^{\mathrm{M}}$ in parallel for an input prompt $\mathbf{x}$}, and (2) ensure that distinct global forking tokens can uniquely initiate these distinct traces.
@@
To do this, we make a simple change to the NTP loss, which now has two requirements: (1) Permutation-invariance: It should not depend on the order of elements in $\mathbf{R}$} and $\bm{g}$}, so we don’t penalize a trace that incurs high NTP loss under one global forking token if the model predicts it well when conditioned on another. (2)No shared global forking token: We want $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ to uniquely initiate distinct reasoning traces, so this requirement prevents conditioning on the same ${\textnormal{g}}^{(i)}$ when generating distinct traces given a question.
@@
To satisfy these requirements, we formally incorporate a subproblem in language modeling: finding the minimum cost bipartite matching configuration between the left vertices $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ and the right vertices $\{{\mathbf{r}}^{(j)}\}_{j=1}^{\mathrm{M}}$ where the cost of each edge between a left vertex $i$ and a right vertex $j$, is the NTP loss of ${\mathbf{r}}^{(j)}$ conditioned on ${\textnormal{g}}^{(i)}$ and an input prompt $\mathbf{x}$}. A matching configuration is a set of edges connecting the left and right vertices where no two edges share a common vertex. Without loss of generality, we assume $\mathrm{N}\geq\mathrm{M}$ to simplify our notation. The total cost involves all vertices on the smaller side of the bipartite graph, so this allows us to write the summation from 1 to $\min\{\mathrm{N},\mathrm{M}\}$, which equals $\mathrm{M}$ under this assumption. We denote a matching configuration as a finite map $\bm{\sigma}:\{1,...,\mathrm{M}\}\rightarrow\{1,...,\mathrm{N}\}$ such that $\bm{\sigma}(j)=i\iff{\mathbf{r}}^{(j)}\text{ is paired with }{\textnormal{g}}^{(i)}$. Let $\bm{\mathfrak{S}}_{\mathrm{P}}:=\{\bm{\sigma}_{k}\}_{k=1}^{\mathrm{P}}$ denote all the $\binom{\mathrm{N}}{\mathrm{M}}\times\mathrm{M}!$ configurations of a bipartite graph. The total cost of each configuration represents the compatibility between $\{{\mathbf{r}}^{(j)}\}_{j=1}^{\mathrm{M}}$ and $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ under this unique matching. Figure 2 shows an example of a matching configuration.
@@
Figure 2: An illustration of one SSFT training step. Step 1: We first construct the cost matrix by evaluating all pairwise combinations: for each ${\mathbf{r}}^{(j)}\in\{{\mathbf{r}}^{(1)},{\mathbf{r}}^{(2)},{\mathbf{r}}^{(3)},{\mathbf{r}}^{(4)}\}$ and each ${\textnormal{g}}^{(i)}\in\{{\textnormal{g}}^{(1)},{\textnormal{g}}^{(2)},{\textnormal{g}}^{(3)},{\textnormal{g}}^{(4)},{\textnormal{g}}^{(5)},{\textnormal{g}}^{(6)}\}$, we compute the NTP loss of ${\mathbf{r}}^{(j)}$ conditioned on ${\textnormal{g}}^{(i)}$ (Equation (4)). Then we use Hungarian algorithm to find $\hat{\bm{\sigma}}$ that minimizes the total bipartite matching cost. Here, this minimum is the sum of the losses highlighted in blue, which means $\hat{\bm{\sigma}}=\{({\textnormal{g}}^{(6)},{\mathbf{r}}^{(1)}),({\textnormal{g}}^{(5)},{\mathbf{r}}^{(2)}),({\textnormal{g}}^{(2)},{\mathbf{r}}^{(3)}),({\textnormal{g}}^{(3)},{\mathbf{r}}^{(4)})\}$. Step 2: We optimize $\bm{\theta}$} by backpropagating the set of NTP losses for ${\mathbf{r}}^{(j)}$, each conditioned on ${\textnormal{g}}^{(\hat{\bm{\sigma}}(j))}$. This is the Hungarian loss in Equation 5.
@@ 2.2 SSFT: Minimizing Set of NTP Losses under Optimal Bipartite Matching
For the first step, we compute all the entries in the cost matrix such as the one in Figure 2 and then apply the Hungarian algorithm (hungarian) to efficiently find the optimal $\hat{\bm{\sigma}}$:
@@
As noted in Equation 4, each matching cost in Equation 3 is the negative log-likelihood of ${\mathbf{r}}^{(j)}$ conditioned on ${\textnormal{g}}^{(\bm{\sigma}(j))}$ under the current model parameters. Here, $\text{sg}(\cdot)$ is stop-gradient, as the matching process is done by discrete optimization w.r.t. $\bm{\sigma}$, so we can save VRAM by not storing intermediate activations. We explicitly indicate that length normalization is done to remove biases toward trace length, so that the matching is driven by semantic content.
After solving $\hat{\bm{\sigma}}$ for each $({\mathbf{x}},{\mathbf{R}})$, our second step optimizes model parameters $\bm{\theta}$} by backpropagating the Hungarian loss in Equation 5. In Algorithm 1, the expectation is replaced by its sample mean over pairs of $({\mathbf{x}},\{{\mathbf{r}}^{(j)}\}_{j=1}^{\mathrm{M}})$ in a mini-batch. In practice, we may use only the first $\mathrm{L}<{\mathrm{T}_{{\mathbf{r}}}}$ tokens in Equation 4 to compute the matching cost and find $\hat{\bm{\sigma}}$ when the training dataset is such that these early tokens reveal sufficient differences. However, we always optimize the Hungarian loss for the full $\mathrm{T}_{{\mathbf{r}}}$} length.
@@
Remarks. The resulting model is not the same as a simple routing of the models independently trained with the nonoverlapping subsets of these traces. Firstly, SSFT allows positive transfer in representation learning within $\{{\mathbf{r}}^{(j)}\}_{j=1}^{\mathrm{M}}$ even though they are matched to different $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$. Secondly, it is not optimal to distill reasoning traces from the same fixed sources for every question if the goal is to maximize both diversity and correctness. Some traces can be filtered by rejection sampling. Our algorithm supports a variable number of target reasoning traces across training steps, with sources that may also change for each $\mathbf{x}$}. Thirdly, even if the two sets of traces, $\{{\mathbf{r}}^{(j)}_{a}\}_{j=1}^{\mathrm{M}}$ for ${\mathbf{x}}_{a}$ and $\{{\mathbf{r}}^{(j)}_{b}\}_{j=1}^{\mathrm{M}}$ for ${\mathbf{x}}_{b}$, are from the same sources, their optimal configurations $\hat{\bm{\sigma}}_{a}$ and $\hat{\bm{\sigma}}_{b}$ can still vary because a teacher model can reason differently under different questions. Lastly, we reserve more global forking tokens than the maximum number of traces ($\mathrm{N}>\mathrm{M}$), and empirically observe that all the forking tokens are being matched throughout the process. This is because the extra forking tokens can maximally intra-differentiate similar traces.

### v3

@@ Abstract
Although LLMs have demonstrated improved performance by scaling parallel test-time compute, doing so relies on generating reasoning paths that are both diverse and accurate. For challenging problems, the forking tokens that trigger diverse yet correct reasoning modes are typically deep in the sampling tree. Consequently, common strategies to encourage diversity, such as temperature scaling, encounter a worsened trade-off between diversity and accuracy. Motivated by this challenge, we treat parallel reasoning as a set-of-next-token-prediction problem and incorporate a set-based global loss into Supervised Fine-Tuning (SFT) using bipartite matching between global forking tokens and unique reasoning traces. We observe that whereas naive fine-tuning with multiple reasoning traces collapses these unique reasoning modes, our proposed method, Set Supervised Fine-Tuning (SSFT), preserves these modes and produces emergent global forking tokens. Global Forking Policy Optimization (GFPO) leverages these maximally steerable tokens to incentivize complex reasoning, and the resulting models consistently outperform their SFT counterparts with GRPO on both math reasoning and execution-based code generation benchmarks.
@@ 1 Introduction
Large language models have recently improved reasoning by allocating more test-time compute to generate more tokens before producing the final answer (OpenAI, 2025). However, extended sequential scaling can lead to “overthinking”, where performance decreases beyond a certain sequence length (Ghosal et al., 2025; Chen et al., 2024a). To mitigate this, another scaling dimension based on repeated parallel sampling and aggregation (Wang et al., 2022; Brown et al., 2024) has shown success in further boosting reasoning performance. However, these methods rely on LLMs generating diverse yet correct solutions; as tasks become harder, a mechanism for increasing diversity is required. Recent work shows that only a minority of tokens in Chain-of-Thought reasoning (Wei et al., 2022) can act as forking tokens that lead to distinct reasoning modes (Wang et al., 2025b), so as the problem becomes harder and the generation becomes longer, it can become substantially harder to sample them. Also, common practices to encourage diversity, typically through temperature scaling, inherently entail a diversity-accuracy trade-off, as the forking tokens that trigger diverse yet correct reasoning modes are typically located deeply within the sampling tree. Moreover, recent theoretical work also shows that increasing the temperature alone does not necessarily guarantee greater diversity unless the model is explicitly trained to ensure coverage. (Verine et al., 2025).
@@
Building on these observations, we aim to leverage diverse reasoning traces to train for coverage (Guo et al., 2025; Google, 2025b). We introduced global forking tokens prior to generate parallel reasoning traces and frame parallel reasoning as a set prediction problem. Specifically, given a question, an LLM, conditioned on a reserved set of tokens in a chosen ordering, generates $\mathrm{M}$ reasoning sequences in parallel, each aligned with one of $\mathrm{M}$ ground-truth reasoning traces. For each ordering, we compute the total autoregressive loss across the generated sequences. By enumerating all possible orderings, we identify the minimum loss, which defines the set language modeling loss, conditioned on the distinct forking tokens (Equation 3). This formulation naturally incorporates coverage into the training objective and is capable of learning global forking tokens that can serve as prompts to trigger reasoning modes that are both diverse and accurate. We operationalize this idea through our Set Supervised Fine-Tuning (SSFT) framework. Our main contributions are summarized below.
@@
  • We introduce global forking tokens and incorporate a set-based loss into SFT via bipartite matching between reserved control tokens and diverse traces (Section 2, Figure 1). The SSFT fine-tuned model globally steer reasoning modes from a single global forking token, reducing dependence on sampling forking tokens mid-generation (Section 2.4; Figure 2).
  • We show that SFT on diverse reasoning traces collapses control tokens into a single reasoning mode (Figure 5), whereas the global forking tokens learned by SSFT initiate distinct reasoning. This is visualized quantitatively through discrepant reasoning effort and accuracy over different control tokens (Figure 4) and improved $\mathrm{Pass@}k$ (Figure 3), and qualitatively through distinct math reasoning strategies across forking tokens (Section A.10).
  • We demonstrate that, across multiple reasoning benchmarks, Global Forking Policy Optimization can leverage the learned global forking tokens to incentivize diverse complex reasoning without collapsing them, and the resulting models consistently outperform their SFT counterparts with GRPO, improving $\mathrm{Pass@}1$ and $\mathrm{Cons@}k$ (Section 3, Table 1, 2).
@@ 2.1 Parallel Reasoning as Set of Next Token Prediction
Problem Setup. In this paper, our goal is not only to instill new reasoning capabilities into a model, but also to ensure that prompting with a set of reserved special tokens, in parallel with a question, elicits distinct reasoning traces. We call these global forking tokens ${\bm{g}}:=\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ instantiated as $\{\texttt{<think\,i>\,}\}_{i=1}^{\mathrm{N}}$ tags. We use ${\textnormal{g}}^{(i)}$ interchangeably with <think i>, depending on context for clarity. We consider a setting with multiple sources of reasoning traces, obtained at low cost without human annotation by distilling from diverse teachers, sampling repeatedly, and potentially filtering with a verifiable metric such as correctness. We adopt this low-cost regime to highlight the effectiveness of our algorithm, though the method extends naturally to settings with well-annotated, human-labeled data. So our problem is to (1) learn to do a set of next-token predictions on multiple distinct yet correct reasoning traces ${\mathbf{R}}:=\{{\mathbf{r}}^{(j)}\}_{j=1}^{\mathrm{M}}$ in parallel for an input prompt $\mathbf{x}$}, and (2) ensure that distinct global forking tokens can uniquely initiate these distinct traces.
@@
To do this, we make a simple change to the NTP loss, which now has two requirements: (1) Permutation-invariance: It should not depend on the order of elements in $\mathbf{R}$} and $\bm{g}$}, so we don’t penalize a trace that incurs high NTP loss under one forking token if the model predicts it well when conditioned on another. (2)No shared global forking token: We want $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ to uniquely initiate distinct reasoning traces, so this requirement prevents conditioning on the same ${\textnormal{g}}^{(i)}$ when generating distinct traces given a question.
@@
To satisfy these requirements, we incorporate a subproblem in language modeling: finding the minimum cost bipartite matching configuration between the left vertices $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ and the right vertices $\{{\mathbf{r}}^{(j)}\}_{j=1}^{\mathrm{M}}$ where the cost of each edge between a left vertex $i$ and a right vertex $j$, is the NTP loss of ${\mathbf{r}}^{(j)}$ conditioned on ${\textnormal{g}}^{(i)}$ and an input prompt $\mathbf{x}$}. A matching configuration is a set of edges connecting the left and right vertices where no two edges share a common vertex. Without loss of generality, we assume $\mathrm{N}\geq\mathrm{M}$ to simplify our notation. The total cost involves all vertices on the smaller side of the bipartite graph, so this allows us to write the summation from 1 to $\min\{\mathrm{N},\mathrm{M}\}$, which equals $\mathrm{M}$ under this assumption. We denote a matching configuration as a finite map $\bm{\sigma}:\{1,...,\mathrm{M}\}\rightarrow\{1,...,\mathrm{N}\}$ such that $\bm{\sigma}(j)=i\iff{\mathbf{r}}^{(j)}\text{ is paired with }{\textnormal{g}}^{(i)}$. Let $\bm{\mathfrak{S}}_{\mathrm{P}}:=\{\bm{\sigma}_{k}\}_{k=1}^{\mathrm{P}}$ denote all the $\binom{\mathrm{N}}{\mathrm{M}}\times\mathrm{M}!$ configurations of a bipartite graph. The total cost of each configuration represents the compatibility between $\{{\mathbf{r}}^{(j)}\}_{j=1}^{\mathrm{M}}$ and $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ under this unique matching. Figure 1 shows an example of a matching configuration.
@@
Figure 1: An illustration of one SSFT training step. Step 1: We first construct the cost matrix by evaluating all pairwise combinations: for each ${\mathbf{r}}^{(j)}\in\{{\mathbf{r}}^{(1)},{\mathbf{r}}^{(2)},{\mathbf{r}}^{(3)},{\mathbf{r}}^{(4)}\}$ and each ${\textnormal{g}}^{(i)}\in\{{\textnormal{g}}^{(1)},{\textnormal{g}}^{(2)},{\textnormal{g}}^{(3)},{\textnormal{g}}^{(4)},{\textnormal{g}}^{(5)},{\textnormal{g}}^{(6)}\}$, we compute the NTP loss of ${\mathbf{r}}^{(j)}$ conditioned on ${\textnormal{g}}^{(i)}$ (Equation (2)). Then we use Hungarian algorithm to find $\hat{\bm{\sigma}}$ that minimizes the total bipartite matching cost. Here, this minimum is the sum of the losses highlighted in blue, which means $\hat{\bm{\sigma}}=\{({\textnormal{g}}^{(6)},{\mathbf{r}}^{(1)}),({\textnormal{g}}^{(5)},{\mathbf{r}}^{(2)}),({\textnormal{g}}^{(2)},{\mathbf{r}}^{(3)}),({\textnormal{g}}^{3},{\mathbf{r}}^{(4)})\}$. Step 2: We optimize $\bm{\theta}$} by backpropagating the set of NTP losses for ${\mathbf{r}}^{(j)}$, each conditioned on ${\textnormal{g}}^{(\hat{\bm{\sigma}}(j))}$.
@@ 2.2 SSFT: Minimizing Set of NTP Losses under Optimal Bipartite Matching
For the first step, we first compute all the entries in the cost matrix such as the one in Figure 1 and then apply the Hungarian algorithm (Kuhn, 1955) to efficiently find the optimal $\hat{\bm{\sigma}}$:
@@
As noted in Equation 2, each matching cost in Equation 1 is the negative log-likelihood of ${\mathbf{r}}^{(j)}$ conditioned on ${\textnormal{g}}^{(\bm{\sigma}(j))}$ under the current model parameters. Here, $\text{sg}(\cdot)$ is stop-gradient, as the matching process is done by discrete optimization w.r.t. $\bm{\sigma}$, so we can save VRAM by not storing intermediate activations. We explicitly indicate that length normalization is done to remove biases toward trace length, so that the matching is driven by semantic content. After solving $\hat{\bm{\sigma}}$ for each $({\mathbf{x}},{\mathbf{R}})$, our second step optimizes model parameters $\bm{\theta}$} by backpropagating on the matching loss in Equation 3. The expectation is replaced by its sample mean over $({\mathbf{x}},\{{\mathbf{r}}^{(j)}\}_{j=1}^{\mathrm{M}})$ in a mini-batch.
Time complexity. In practice, we use only the first $\mathrm{L}<{\mathrm{T}_{{\mathbf{r}}}}$ tokens in Equation 2 to compute the matching cost when early tokens already reveal sufficient differences between traces. This $k=\frac{{\mathrm{T}_{{\mathbf{r}}}}}{\mathrm{L}}$ factor reduces the time complexity of matching to $\mathcal{O}(\frac{NM}{k})$. Backpropagation is performed only on the $\mathrm{M}$ matched sequences and can use the full $\mathrm{T}_{{\mathbf{r}}}$} length at the same compute by SFT.
@@
Remarks. The resulting model is not the same as a simple routing of the models independently trained with the nonoverlapping subsets of these traces. Firstly, SSFT allows positive transfer in representation learning within $\{{\mathbf{r}}^{(j)}\}_{j=1}^{\mathrm{M}}$ even though they are matched to different $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$. Secondly, it is not optimal to distill reasoning traces from the same fixed sources for every question if the goal is to maximize both diversity and correctness. Our algorithm supports a variable number of target reasoning traces across training steps, with sources that may also change for each $\mathbf{x}$}. Thirdly, even if the two sets of traces, $\{{\mathbf{r}}^{(j)}_{a}\}_{j=1}^{\mathrm{M}}$ for ${\mathbf{x}}_{a}$ and $\{{\mathbf{r}}^{(j)}_{b}\}_{j=1}^{\mathrm{M}}$ for ${\mathbf{x}}_{b}$, are from the same sources, their optimal configurations $\hat{\bm{\sigma}}_{a}$ and $\hat{\bm{\sigma}}_{b}$ can still vary because a teacher model can reason differently under different questions. Lastly, we reserve more global forking tokens than the maximum number of traces ($\mathrm{N}>\mathrm{M}$), and empirically observe that all the forking tokens are being matched throughout the process. This is because the extra forking tokens can maximally intra-differentiate similar traces.
2.3 GFPO: Optimizing the Selection of Global Forking Token
Although SSFT can leverage unbalanced bipartite graphs to learn correlations between questions and forking tokens given sufficient data, collecting additional traces is often unnecessarily expensive once the model already learns distinct reasoning patterns. Instead, we apply a small number of RL steps using only the policy gradients of global forking tokens—an extremely efficient variant of Wang et al. (2025b; a). This Global Forking Policy Optimization (GFPO) properly shapes the output distribution of ${\textnormal{g}}^{(i)}$ conditioned on $\mathbf{x}$}. It can be implemented trivially by adding a few lines of Python slicing to any off-the-shelf code (Sheng et al., 2024), since global forking tokens always begin at the same indices of generations. Full generations are used solely to compute the advantage of ${\textnormal{g}}^{(i)}$, and they do not participate in backpropagation.